OpenStudy (anonymous):

expand and simplify x(x-3) - (3-2x)^2

OpenStudy (anonymous):

x(x-3) - (3-2x)^2 =x^2 - 3x - (9 - 12x + 4x^2) =x^2 -3x - 9 + 12x - 4x^2 = - 3x^2 + 9x - 9 Dividing by - 3 = x^2 - 3x + 3 ANS